OpenStudy (anonymous):

1/27=9^8x-4 did i get it right? x=0.114140?

OpenStudy (dumbcow):

-3 = 16x -8 5 = 16x 5/16 = x

OpenStudy (anonymous):

3^(-3) = 3^(16x-8) -3 = 16x - 8 5 = 16x x = 5/16 or 0.3125

OpenStudy (anonymous):

i cant get these and i dont know y